OpenStudy (pphalke):

the numbers on two consecutively numbered gym lockers have a sum of 135. what are the locker numbers?

OpenStudy (welshfella):

n + (n + 1) = 135 solve for n

OpenStudy (pphalke):

how do i do that

OpenStudy (photon336):

Consecutive means that each term is increased by one. so two consecutive numbers, just a fancy way of saying 1 more than. the first number would be n the second would be n+1 these two are consecutive because they differ by 1. so the sum n+(n+1) = 135

OpenStudy (photon336):

@pphalke

OpenStudy (pphalke):

how do i solve this @Photon336

OpenStudy (photon336):

can you try to do it first then ill guide you

OpenStudy (photon336):

you need to combine like terms. n+n+1 = 135

OpenStudy (pphalke):

2n+1=135

OpenStudy (pphalke):

then i subtract from 1 and get 134

OpenStudy (welshfella):

ok now subtract 1 from both sides

OpenStudy (welshfella):

yea now divide both sides by 2

OpenStudy (pphalke):

n equals 67

OpenStudy (welshfella):

correct

OpenStudy (pphalke):

the locker numbers are 67 and 68

OpenStudy (welshfella):

so thats the first locker number . What's the one next to it?

OpenStudy (welshfella):

yes